A boat having a length of 3 metre and breadth 2 metre is floating on a lake . The boat sinks by 1 cm when many sinks on it. What is the mass of the man?

Grade:9

1 Answers

Arun
6 years ago
first we will calculate the volume of water displaces then will multiply with the density of water.
Volume of water displaced = 3*2*0.01 = 0.06 m cube

Mass of Man = Volume of water displaced * Density of water

= 0.06 * 1000 = 60 kg